2 machines produce washers. The 1st machine can produce twice as many washers in one hour than the 2nd machine. If the 1st machine runs 2 hours and the 2nd machine runs 3 hours and 560 washers are produced, how many does each one produce in one hour?
Answer by ewatrrr(24785)


Hi,
Let x represent the number Of washers the second machine can produce in one hour
Then 2x is the number Of washers the first machine can produce in one hour
Write as we Read
2hr *2x/hr + 3hr*x/hr = 560 washers
solve for x
4x + 3x = 560
7x = 560
x = 80, number Of washers the second machine can produce in one hour
(2*80) Or 160 is number Of washers the first machine can produce in one hour
checking our answer
2*160 + 3*80 = 560